A jetplane and a tanker that are 525 mi apart head toward each other so that the jet can refuel. The jet flies 200 mi/h faster than the tanker. Determine the speed of each aircraft if they meet in 45 minutes?
Answer by oberobic(2304)

Keep in mind d=rt (distance = rate times time).
We know the two aircraft are 525 miles from each other at the start.
We assume they fly in a straight line toward each other to be the shortest distance.
We know both fly for 45 minutes = 3/4 of an hr to come together. (The speeds are in miles per hour, so you have to get the time in hours, not minutes.)
And we know that one plane flies at x mph and the other flies at x+200 mph.
The first plane flies d1 = x*3/4 = 3/4x
The second plane flies d2 = (x+200)*3/4 = 3/4x +150
So, d2 = d1 + 150.
Given d1 + d2 = 525, then d1 + d1 + 150 = 525. Subtracting 150 from both sides,
2d1 = 375.
Dividing by 2.
d1 = 187.5
d2 = 187.5 + 150 = 337.5
Looking back at what we know, the first plane flies d = rt. 187.5 = 3/4x.
Dividing by 3/4 (or .75 decimal value), we find x = 187.5/.75 = 250 mph.
The second plane flies x+200 mph = 250 + 200 = 450 mph.
ALWAYS check!
d1 = 3/4*250 = 187.5
d2 = 3/4*450 = 337.5
187.5 + 337.5 = 525